# MixDQ: Memory-Efficient Few-Step Text-to-Image Diffusion Models with Metric-Decoupled Mixed Precision Quantization

### v1

@@ 3.2 Metric-Decoupled Sensitivity Analysis
We look into this issue and observe the following phenomena in Fig. 5 (Right): (1) Both changes in image content and quality impact the SQNR. In particular, compared with the 30.03dB image, the SQNR decline of an image with changed contents but high quality (-3.51dB) is much more significant than that of an image with similar contents but unacceptable quality (-0.26dB). This means that when evaluating images with varying content and quality, SQNR tends to underestimate the performance degradation resulting from decreased image quality. (2) Image quality and content are mainly affected by different layer types. Specifically, quantizing the cross-attention layers causes content changes, but the quality remains good. This aligns with the intuition that text supervision predominantly interacts with cross-attention layers, which control the image content. Conversely, when quantizing convolutions, the quality degrades and the content is preserved. Further details of the correlation between layer types and their impact on content or quality are provided Sec. 0.D of the appendix.
@@ 4.2 Performance and Efficiency Comparison
We experiment with two quantization settings: the “weight-only” scheme which seeks larger compression rate of model size, and the “normal” scheme focuses on both storage saving and latency speedup. During activation quantization, we observed that certain layers retain sensitivity and cannot be quantized to 8 bits without sacrificing performance. To address this, we choose to retain 1% of the most sensitive layers based on metric-decoupled sensitivity. Appendix 0.C shows that it introduces minimal overhead but ensures the preservation of performance.
@@ 4.3 Hardware Resource Savings
Memory footprint reduction. Fig. 0.C (a) shows the GPU memory usage of MixDQ and the FP16 baseline on UNet inference. MixDQ can reduce the memory footprint from two aspects. Firstly, quantizing the model weights leads to smaller allocated memory to store the UNet model. Secondly, quantizing the activations saves allocated memory to store residual connection activations. Combining the two benefits, we can effectively reduce the peak memory footprint by 1.87$\times$, 3.03$\times$ and 3.03$\times$ under the W8A8, W4A16 and W4A8, respectively.
Speedup of model inference. Fig. 0.C (a) shows the latency of the UNet model inference on Nvidia RTX 4080. We use the W8A8 scenario to showcase the speedup. MixDQ W8A8 can accelerate the inference of UNet by 1.52$\times$ over the FP16 baseline. 0.C (b) gives a breakdown of the inference latency. There are three types of layers in the model: quantizable layers, including linear and convolution layers; non-quantizable layers, such as normalization and non-linear activation layers; and quantization layers, that perform the FP16-to-int8 conversion. The non-quantizable layers have the same latency in both the baseline and MixDQ. The quantizable layers are accelerated by 1.97$\times$, approximately the same ratio between INT8 and FP16 hardware peak throughput on RTX4080 (2$\times$). MixDQ requires quantization layers for converting activations, but even with this overhead, the end-to-end acceleration still reaches 1.52$\times$.
@@ Appendix 0.C Detailed Analysis of Hardware Experiments
(1) Illustrated in Fig. 12(b), the W8A8 bar displays an almost negligible red segment, corresponding to the "1% most sensitive activation layers retained from FP16" as discussed in Sec. 4.2 of the main paper. This component introduces minimal overhead while significantly contributing to the preservation of image quality.
(2) Observing the table in Appendix 0.C, we note that W4A16 quantization solely achieves memory savings, while the latency remains consistent with the FP16 baseline. The reduction in memory primarily stems from the decreased size of the model parameters, which are transferred to the GPU and serve as the "static memory cost." However, the W4A16 model continues to utilize FP16 computation, thus maintaining similar latency.
(3) Observing the table in Appendix 0.C, W4A16 and W4A8 exhibit similar memory optimization. The key distinction lies in the fact that W4A8 quantizes the activations, thereby reducing the memory cost through saved activations, primarily affecting the shortcut feature for U-Net. However, we discover that the occupied size by these activations is relatively small, amounting to less than 100MB compared to the model size of 6GB. Consequently, the additional savings attributed to activation quantization are not evident.
(4) Observing the table in Appendix 0.C, both W4A8 and W8A8 demonstrate a noteworthy 1.52× latency speedup compared to the FP16 baseline. This acceleration is attributed to the utilization of INT8 GPU kernels. It’s worth noting that we haven’t implemented the INT4 operators, making W4A8 and W8A8 quite similar in terms of latency improvement. However, W4A8 has the potential to achieve even higher latency speedup.
Bit-width Memory Latency FID ($\downarrow$)
(W/A) Opt. Opt.
16/16 - - 17.15
4/16 3.03$\times$ - 17.23
8/8 1.87$\times$ 1.52$\times$ 17.03
4/8 3.03$\times$ 1.52$\times$ 17.68

### v2

@@ 3.2 Metric-Decoupled Sensitivity Analysis
We look into this issue and observe the following phenomena in Fig. 5 (Right): (1) Both changes in image content and quality impact the SQNR. In particular, compared with the 30.03dB image, the SQNR decline of an image with changed contents but high quality (-3.51dB) is much more significant than that of an image with similar contents but unacceptable quality (-0.26dB). This means that when evaluating images with varying content and quality, SQNR tends to underestimate the performance degradation resulting from decreased image quality. (2) Image quality and content are mainly affected by different layer types. Specifically, quantizing the cross-attention layers causes content changes, but the quality remains good. This aligns with the intuition that text supervision predominantly interacts with cross-attention layers, which control the image content. Conversely, when quantizing convolutions, the quality degrades and the content is preserved. Further details of the correlation between layer types and their impact on content or quality are provided in Appendix 0.D.
@@ 4.1 Experimental Settings
Table 2: Comparison with other quantization techniques for W8A8 on COCO annotations subset. The “CLIP” and “IR” denotes CLIP Score and ImageReward metric. The metrics are evaluated with the first 1024 prompts in COCO annotations. The “MP-only” represents adopting the mixed precision only.
Model Method FID($\downarrow$) CLIP($\uparrow$) IR($\uparrow$)
SDXL-turbo (1 step) FP16 84.51 0.26 0.84
Naive PTQ 165.92 (+81.4) 0.15 (-0.11) -1.72 (-2.56)
PTQD [5] 340.74 (+256.2) 0.12 (-0.14) -2.28 (-3.12)
Q-Diffusion [13] 149.15 (+64.6) 0.16 (-0.10) -1.69 (-2.53)
EDA-DM [16] 137.98 (+53.5) 0.16 (-0.10) -1.71 (-2.55)
MP-only [54] 114.70 (+30.2) 0.15 (-0.11) -0.61 (-1.45)
Non-Uniform (FP8) [23] 101.73 (+17.2) 0.24 (-0.01) 0.16 (-1.45)
MixDQ 83.39 (-1.12) 0.27 (+0.01) 0.84 (+0.00)
@@ 4.2 Performance and Efficiency Comparison
We experiment with two quantization settings: the “weight-only” scheme which seeks larger compression rate of model size, and the “normal” scheme focuses on both storage saving and latency speedup. During activation quantization, we observed that certain layers retain sensitivity and cannot be quantized to 8 bits without sacrificing performance. To address this, we choose to retain 1% of the most sensitive layers based on metric-decoupled sensitivity. Fig. 7 shows that it introduces minimal overhead but ensures the preservation of performance.
@@
We further compare MixDQ’s performance with other quantization techniques for W8A8 in Tab. 2. For PTQD [5], we conduct linear regression to determine $K$. However, we observe a relatively lower linear correlation (0.59) compared to the original paper. In the case of 1-step generation, the sampling process is deterministic, the "uncorrelated noise correction" inapplicable, and only the "correlated noise correction" is adopted. Despite being effective for multi-step quantization, we observe that PTQ causes notable quality degradation for the challenging 1-step quantization. EDA-DM [16] improves upon the quantization parameter tuning in Q-Diffusion [13] with layer-wise reconstruction. However, only a marginal improvement is witnessed. The term "MP-only" refers to "mixed precision only" with the SQNR-based method [54], where due to suboptimal quantization sensitivity analysis, a significant performance degradation is still evident. The "Non-uniform" refers to the use of the FP8 (Floating-point 8-bit) format [23], which introduces exponential bits to handle a larger dynamic range of data distribution. While this notably improves performance, it still falls short of the FP16 baseline. In contrast, MixDQ achieves W8A8 quantization without any performance loss.
@@ 4.3 Hardware Resource Savings
Memory footprint reduction. Fig. 7 (a) shows the GPU memory usage of MixDQ and the FP16 baseline on UNet inference. MixDQ can reduce the memory footprint from two aspects. Firstly, quantizing the model weights leads to smaller allocated memory to store the UNet model. Secondly, quantizing the activations saves allocated memory to store residual connection activations. Combining the two benefits, we can effectively reduce the peak memory footprint by 1.87$\times$, 3.03$\times$ and 3.03$\times$ under the W8A8, W4A16 and W4A8, respectively.
Speedup of model inference. Fig. 7 (a) shows the latency of the UNet model inference on Nvidia RTX 4080. We use the W8A8 scenario to showcase the speedup. MixDQ W8A8 can accelerate the inference of UNet by 1.52$\times$ over the FP16 baseline. 7 (b) gives a breakdown of the inference latency. There are three types of layers in the model: quantizable layers, including linear and convolution layers; non-quantizable layers, such as normalization and non-linear activation layers; and quantization layers, that perform the FP16-to-int8 conversion. The non-quantizable layers have the same latency in both the baseline and MixDQ. The quantizable layers are accelerated by 1.97$\times$, approximately the same ratio between INT8 and FP16 hardware peak throughput on RTX4080 (2$\times$). MixDQ requires quantization layers for converting activations, but even with this overhead, the end-to-end acceleration still reaches 1.52$\times$.
@@ 5.1 Analysis of Paretor Frontier
5.2 Analysis of BOS-aware Quantization
We further conduct experiments to demonstrate the effectiveness of BOS-aware Quantization. As discussed in Section 3.1, the first BOS token in CLIP embedding exhibits significantly larger values compared to others. While some specialized quantization techniques, such as non-uniform quantization, could address this outlier issue, introducing FP8 quantization for text embedding still incurs notable performance degradation (See Fig. 9). We also visualize the distribution of the BOS token and observe large variance along the channel dimension. Certain outlier channels have values 800 $\times$ greater than other channels, as depicted in Fig. 10. Additionally, Fig. 10 illustrates that in the T5 [11] text encoder adopted by more recent papers [1, 50], although the first token is no longer an outlier, the channel imbalance problem persists. It’s important to highlight that BOS-aware quantization suggests that the quantization of the fixed BOS feature could be skipped, resolving all its quantization issues beyond mere "outliers".
Figure 9: The comparison of different quantization techniques for text embedding. The content preservation notably improves for FP8 quantization, however, is still inferior to BOS-aware
Figure 10: The visualization of distribution for T5/CLIP text embeddings. Left: the histogram for BOS and other tokens. Right: the channel-wise distribution of the BOS token.
@@ 6 Conclusions
Acknowledgements
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist). We thank for all the support from Infinigence-AI.
@@ Appendix 0.C Detailed Analysis of Hardware Experiments
(1) Illustrated in Fig. 7(b) in the main paper Sec. 4.3, the W8A8 bar displays an almost negligible red segment, corresponding to the "1% most sensitive activation layers retained from FP16" as discussed in Sec. 4.2 of the main paper. This component introduces minimal overhead while significantly contributing to the preservation of image quality.
(2) Observing the table in Fig. 7 in the main paper Sec. 4.3, we note that W4A16 quantization solely achieves memory savings, while the latency remains consistent with the FP16 baseline. The reduction in memory primarily stems from the decreased size of the model parameters, which are transferred to the GPU and serve as the "static memory cost." However, the W4A16 model continues to utilize FP16 computation, thus maintaining similar latency.
(3) Observing the table in Fig. 7 of the main paper Sec. 4.3, W4A16 and W4A8 exhibit similar memory optimization. The key distinction lies in the fact that W4A8 quantizes the activations, thereby reducing the memory cost through saved activations, primarily affecting the shortcut feature for U-Net. However, we discover that the occupied size by these activations is relatively small, amounting to less than 100MB compared to the model size of 6GB. Consequently, the additional savings attributed to activation quantization are not evident.
(4) Observing the table in Fig. 7 of the main paper Sec. 4.3, both W4A8 and W8A8 demonstrate a noteworthy 1.52× latency speedup compared to the FP16 baseline. This acceleration is attributed to the utilization of INT8 GPU kernels. It’s worth noting that we haven’t implemented the INT4 operators, making W4A8 and W8A8 quite similar in terms of latency improvement. However, W4A8 has the potential to achieve even higher latency speedup.